# Navigating the Impact of Structured Output Format on Large Language Models through the Compass of Causal Inference

### v1

@@ Appendix F Small language models
Figure 19: Successful response rate of different LLMs under zero-shot and few-shot Prompts.
@@ B Standard & JSON format
Appendix G Additional interventions
@@
Figure 22: Interventions of context description and problem formulation on the first sample from GMS8K.
@@ C LLMs’ generation JSON format - - -
Appendix H Multiple trials
@@
Appendix I ELLC benchmark on complex tasks
We extended the LLC dataset to develop the Extended Last Letter Concatenation (or ELLC) dataset, which introduces additional complexity in both symbolic and linguistic reasoning. Symbolic reasoning is enhanced by increasing the number of words and requiring extraction from more challenging positions, such as middle letters. Linguistic reasoning adds a further step: forming common English words from the concatenated letters, thus assessing linguistic capabilities.
@@ Appendix I ELLC benchmark on complex tasks
Appendix J Concluding remarks

### v2

@@ Appendix G Small language models
Figure 19: Successful response rate of different SLMs under zero-shot and few-shot Prompts.
@@ B Standard & JSON format
Appendix H Additional interventions
@@
Figure 22: Interventions of context description and problem formulation on the first sample from GSM8K.
@@ C LLMs’ generation JSON format - - -
Appendix I Multiple trials
@@
Appendix J ELLC benchmark on complex tasks
We extended the LLC dataset to develop the Extended Last Letter Concatenation (or ELLC) dataset, which introduces additional complexity in both symbolic and linguistic reasoning via an evolutionary strategy (Li et al., 2024a). Symbolic reasoning is enhanced by increasing the number of words and requiring extraction from more challenging positions, such as middle letters. Linguistic reasoning adds a further step: forming common English words from the concatenated letters, thus assessing linguistic capabilities.
@@
Appendix K Concluding remarks